OpenStudy (anonymous):

Solve y equals two times x plus three for x. x equals y divided by negative two minus three x equals y plus three divided by two x equals y minus three divided by negative two x equals y divided by three plus two

OpenStudy (anonymous):

hi my first time here

OpenStudy (anonymous):

i need help with a answer

OpenStudy (imstuck):

Hello! Is this solving a system of 3 equations simultaneously?

OpenStudy (imstuck):

I'm not sure exactly what it is you are asking for in your post.

OpenStudy (anonymous):

Solve y equals two times x plus three for x. x equals y divided by negative two minus three x equals y plus three divided by two x equals y minus three divided by negative two x equals y divided by three plus two

OpenStudy (imstuck):

y=2x+3. And you need to solve that for x. Ok, the first thing you need to do is move the 3 to the other side of the equals sign. If the 3 is added to the one side, do you know how to move from that side to the other side? What mathematical operation would you use?

OpenStudy (anonymous):

well in my paper i put v=3.14 r^2h

OpenStudy (imstuck):

??????????????????? What is that v? I'm lost now...

OpenStudy (anonymous):

volume

OpenStudy (texaschic101):

you are not dealing with volume in this problem

OpenStudy (anonymous):

oooo sorry im not good with math

OpenStudy (imstuck):

Did you maybe mix up two problems?

OpenStudy (anonymous):

its tell me that y=-2x +3

OpenStudy (imstuck):

The problem you are solving for x is y = 2x + 3. You have to get the x all by itself, with nothing 'attached' to it and nothing on the same side of the = sign. So that means you have to move the 3 to the other side. It's added to the 2x, so in order to move it you have to subtract it, like this: y-3=2x+3-3--> y-3=2x is what you have now.

OpenStudy (texaschic101):

I think the problem is actually y = -2x + 3

OpenStudy (imstuck):

y-3 = 2x. Now you have to move the 2 away from the x. The 2 is multiplied by the x, so in order to "undo" the multiplication, you have to divide. Divide both sides by 2 like this:\[\frac{ y-3 }{ 2 }=\frac{ 2x }{ 2 }\]

OpenStudy (imstuck):

The 2's on the right side will cancel each other out, leaving you only with an x. So your answer to this is:\[\frac{ y-3 }{ -2 }=x\] The third choice down.

OpenStudy (anonymous):

i got it thank alot

OpenStudy (imstuck):

You're welcome!

OpenStudy (anonymous):

so hey only u guys r on